OpenStudy (tinalikescats):

Consider the equations y = |x − 1| and y = 3x + 2. The approximate solution of this system of equations is a.(-0.2, 1.2) b.(0.2, -1.2) c.(0.2, 1.2) d.(-0.2, -1.2)

OpenStudy (lollylau):

From the first equation: y^2 = x^2 - 2x + 1 From the second equation: y^2 = 9x^2 + 12x + 4 Subtracting one from the other: 8x^2 + 14x - 3 = 0 Find x, then substitute x into one of the equations to obtain y.

OpenStudy (tinalikescats):

Idk how to do that

OpenStudy (tinalikescats):

@LollyLau

OpenStudy (lollylau):

Absolute value means the result must be non-negative. By squaring it we remove the absolute value symbol and replace it with a square instead, which is easier to work with. The equation of order two is solved with the quadratic equation.

OpenStudy (tinalikescats):

I forget how to do quadratic equation :( math used to be my #1 subject and I had a seizure so I forgot everything in life

OpenStudy (lollylau):

\[x=\frac{-b \pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (tinalikescats):

._. I don't understand any of that

OpenStudy (lollylau):

Alternatively, you can graph the two equations and read the values of the intersections.

OpenStudy (misty1212):

HI!!

OpenStudy (tinalikescats):

ugh :/ I still don't understand

OpenStudy (misty1212):

ok forget the quadratic equation

OpenStudy (tinalikescats):

alright that's forgotten

OpenStudy (misty1212):

if \(x>1\) then \(|x-1|=x-1\) solve \[x-1=3x+2\] in three steps

OpenStudy (tinalikescats):

I think I solved it backwards cause I got -1.5

OpenStudy (misty1212):

\[x-1=3x+2\\ -1=2x+2\\ -3=2x\\ -\frac{3}{2}=x\] but that is not a solution, because we assumed that \(x>1\)and it isn't so now one more step

OpenStudy (misty1212):

if \(x<1\) then \(|x-1|=1-x\) solve \[1-x=3x+2\]

OpenStudy (misty1212):

let me know what you get

OpenStudy (tinalikescats):

pretty positive I did it totally wrong I got -3.3

OpenStudy (misty1212):

add \(x\) to both sides then subtract 2 from both sides, then divide both sides by 4

OpenStudy (tinalikescats):

-.25?

OpenStudy (misty1212):

yes

OpenStudy (tinalikescats):

awesome :DD

OpenStudy (misty1212):

which for some reason they rounded to \(-.2\)

OpenStudy (tinalikescats):

Someone please help with this last part?